### v1

Abstract
The 8.4 eV transition in the 229Th nucleus is the basis for a high-precision nuclear clock with exceptional sensitivity to new physics effects. We have identified several cases in the Th+ ion where electronic excitations closely resonate with the nuclear excitation, with the smallest energy difference being $\Delta=-0.09$ cm-1. We investigate the electronic bridge process, in which nuclear excitation is induced via electronic transitions, and demonstrate that a proper selection of laser frequencies can lead to a dramatic enhancement of this effect. Additionally, we show that the interaction with electrons significantly shortens the lifetime of the nuclear excited state.
I Introduction
It was suggested in Ref.[1] that the 229Th nucleus has a low-energy metastable state. Ref.[2] proposed using the transition between the ground and metastable states as the basis for a nuclear clock. The precision of such a clock in a ”stretched” state, with maximal projections of both electronic and nuclear angular momenta, is expected to be very high because, in this state, the electron and nuclear variables are not entangled [3]. Indeed, in such a state, the nucleus is practically decoupled from external fields, such as those produced by blackbody radiation. This leads to a significant suppression of corresponding systematic effects.
On the other hand, the frequency of this transition is highly sensitive to effects of physics beyond the Standard Model, including violations of the Lorentz invariance and the Einstein equivalence principle [4], as well as variations in fundamental physical constants [5], such as the fine-structure constant $\alpha$, quark masses, and the QCD scale $\Lambda_{QCD}$. This sensitivity arises from the very small transition energy (8.4 eV), which is the result of strong cancellations between contributions from strong and electromagnetic nuclear interactions, typically at the MeV scale. This cancellation makes the transition frequency extremely sensitive to small changes in these contributions caused by variations in fundamental constants. Such variations may result from interactions with scalar or axion dark matter fields. Therefore, measuring these variations in clock experiments may provide an efficient method for searching for these fields [6, 7].
The measurement of the energy of the nuclear clock transition in 229Th has been an ongoing effort for many years (see, e.g., [8, 9, 10]). Recent advances in laser spectroscopy methods have led to a breakthrough in accuracy[11, 12, 13]. The transition frequency has been measured to be [12] $\omega_{N}=2,020,407,384,335(2)$ kHz [13] (67393 cm-1) in 229Th atoms embedded in a solid. Even higher accuracy is anticipated in future experiments with 229Th ions [3].
The amplitude of the nuclear M1 transition is suppressed by five orders of magnitude compared to typical atomic E1 transitions. However, it is possible to enhance this transition using the electronic bridge (EB) process, where electronic transitions induce nuclear transitions via hyperfine interaction (hfi). The EB effect also increases the probability of decay of the nuclear excited state. Corresponding calculations have been performed for Th II, Th III, and Th IV ions [14, 15, 16, 17, 18, 19, 20, 21].
Previous EB calculations for the decay [15] and excitation [16] of the excited nuclear state in the Th II ion could not provide accurate results due to large errors in the measured value of the nuclear excitation energy and the absence of measurements of the positions of Th II electron energy levels near the nuclear excitation energy. According to theoretical expectations [22], the spacing between electron energy levels in this region is very small, making resonance with the nuclear excitation possible.
The aim of this paper is to propose an atomic excitation scheme that maximizes the EB effect and to estimate the magnitude of this effect. Additionally, we estimate the impact of the EB process on the decay of the nuclear excited state. Recent data on the Th II electron spectrum [23] (see Supplemental Material for the most complete data) allow us to identify several cases where a small energy denominator in the EB excitation formula leads to a very large resonance enhancement of the EB effect.
The error in the calculated level positions exceeds the very small spacing between the measured electron energy levels in highly excited states. To estimate the EB effects, we adopt a statistical approach (based on the calculations of many possible EB transition probabilities) due to the current lack of information needed to match measured energy levels with calculated electron wave functions. Accurate measurements of the magnetic moments and angular momenta of the Th II electron states would assist in the theoretical identification of the measured levels and improve the accuracy of EB predictions.
II Electronic bridge for nuclear excitation and decay in Th II
We follow the method used in Refs. [16, 21] for the Th II anf Th III ions. The decay of any atomic state with energy $E_{n}>\omega_{N}$ may include nuclear excitation. Since the nuclear excitation energy $\omega_{N}=67393$ cm-1 lies outside the optical region, we consider a two-step excitation of the electronic states(see also [16]). In the first step, the atom is excited from the ground state (GS) to one of the intermediate states $t$ used in [16]. In the second step, the ion is further excited by a second laser with frequency which is chosen to satisfy the energy conservation condition for simultaneous nuclear excitation and excitation of the final electronic state, $\omega_{1}+\omega_{2}=\omega_{N}+E_{s}$ ($\omega_{1}=E_{t}$). Nuclear excitation is due to the hyperfine interaction between atomic electrons and the nucleus. Corresponding diagram is presented on Fig. 1. Its calculation involves a summation over all intermediate states $n$. However, in this analysis, we focus on the dominant contributions, which come from specific intermediate states $n$ and final states $s$ providing a very small energy denominator in Eq. (3). Note that the final state $s$ is not the ground state but a low-lying excited state satisfying selection rules for the hfi-induced E1 transition from the state $t$ (see diagram on Fig. 1). We do not go to highly excited states $s$ to keep laser frequencies within optical region. For each final electron state $s$, we look for an intermediate state $n$ (using energy level data from Ref. [23]) which corresponds to the smallest energy denominator $\Delta$ in (3), $\Delta=E_{n}-E_{s}-\omega_{N}$, and use this state for estimation of the EB amplitude.
Figure 1: A diagram for the EB excitation process in Th II. It is assumed that the atom is in initial electronic state $t$; $n$ and $s$ are intermediate and final electronic states, $g$ ground nuclear state, $m$ is isomeric nuclear state.
The rate of an induced excitation from state $b$ to state $a$ $W_{ba}^{\rm in}$ can be calculated using the rate of a spontaneous transition $a\rightarrow b$ $W_{ab}$ [25]:
$W_{ba}^{\rm in}=W_{ab}\frac{4\pi^{3}c^{2}}{\omega^{3}}I_{\omega}.$ (1)
Here $I_{\omega}$ is the intensity of isotropic and unpolarized incident radiation. Index $b$ corresponds to nuclear ground state and electron excited state, index $a$ corresponds to nuclear excited state and electronic ground or low energy excited state. The rate of a spontaneous transition via electronic bridge process is given by [14]
$W_{ab}=\frac{4}{9}\left(\frac{\omega}{c}\right)^{3}\frac{|\langle I_{g}||M_{k}||I_{m}\rangle|^{2}}{(2I_{m}+1)(2J_{t}+1)}G_{2}^{(k)},$ (2)
Keeping in mind the relation (1), we assume that $\omega$ in (2) is the frequency of second excitation ($\omega\equiv\omega_{2}$). Factor $G_{2}$ in (2) depends on electrons only. It corresponds to upper line of Fig. 1. Assuming resonance situation and keeping only one strongly dominating term, we have
$G_{2}^{(k)}\approx\frac{1}{2J_{n}+1}\left[\frac{\langle s||T_{k}||n\rangle\langle n||D||t\rangle}{\omega_{ns}-\omega_{N}}\right]^{2}.$ (3)
Here $T_{k}$ is the electron part of the hyperfine interaction operator (magnetic dipole (M1) for $k=1$ and electric quadrupole (E2) for $k=2$), $D$ is the electric dipole operator (E1). States $n$ and $s$ are chosen to get close to a resonance, $\epsilon_{n}-\epsilon_{s}\equiv\omega_{ns}\approx\omega_{N}$.
It is convenient to present the results in terms of dimensionless ratios ($\beta$) of EB transition rates to the nuclear transition rates.
$\beta_{M1}=\Gamma^{(1)}_{\rm EB}/\Gamma_{N}(M1),\ \ \beta_{E2}=\Gamma^{(2)}_{\rm EB}/\Gamma_{N}(E2).$ (4)
Here $\Gamma^{(k)}$ is given by (2) ($\Gamma^{(k)}\equiv W_{ab}$). Both parameters, $\beta_{M1}$ and $\beta_{E2}$ can be expressed via $G2$, (Eq. (3)) [17]
$\beta_{M1}=\left(\frac{\omega}{\omega_{N}}\right)^{3}\frac{G_{2}^{(1)}}{3(2J_{s}+1)},$ (5)
$\beta_{E2}=\left(\frac{\omega}{\omega_{N}}\right)^{3}\frac{4G_{2}^{(2)}}{k_{n}^{2}(2J_{s}+1)}.$ (6)
The $M1$ and $E2$ contributions can be combined into one effective parameter $\tilde{\beta}$ using the known ratio of the widths of the nuclear $M1$ and $E2$ transitions, $\gamma=\Gamma_{\gamma}(E2)/\Gamma_{\gamma}(M1)=6.9\times 10^{-10}$, [26]. Then $\tilde{\beta}=\beta_{M1}(1+\rho)$, where $\rho=\gamma\beta_{E2}/\beta_{M1}$.
III Numerical calculations of the electronic bridge contribution to the nuclear excitation
Table 1: Versions of the EB process with resonant nuclear excitation. The notations $t$, $n$, and $s$ refer to the diagram on Fig. 1. It is assumed that $E_{t}$ = 36390.53 cm-1, $\omega_{2}=\omega_{N}+E_{s}-E_{t}$ is the frequency of second laser excitation. The energy denominator in Eq. (3) is given by $\Delta=E_{n}-E_{s}-\omega_{N}$. The third column lists all possible values of $J_{n}$ from the experimental data in Ref. [23], the last two columns present the calculated values of the EB enhancement factors $\beta$ for specific $J_{n}$. Here, $x\equiv\langle s||T_{k}||n\rangle^{2}\langle n||D||t\rangle^{2}$, and $\langle x\rangle$ represents the average value of $x$ over theoretical states that are candidates to represent the actual states $t$ and $n$. No averaging is performed over $s$ since it is a low-lying state that is easily identified in the calculations. The median value of $x$ is denoted by $x_{m}$. Numbers in square brackets indicate powers of ten.
$N$ $E_{n}$ [^1] Ref. [23], Supplemental Material. $J_{n}$ $E_{s}$222 NIST database, [24]. $J_{s}$ $\omega_{2}$ $\Delta$ $\beta(\langle x\rangle)$ $\beta(x_{m})$
[cm-1] [cm-1] [cm-1] [cm-1]
1 73637.54 3/2 or 5/2 6244.29 1/2 37247.10 -0.09 5.5[6]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole hfi. 1.6[5]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole hfi.
2 73637.54 3/2 or 5/2 6244.29 1/2 37247.10 -0.09 1.2[4]444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole hfi contributes. 660444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole hfi contributes.
3 74396.40 3/2 or 5/2 7001.42 3/2 38004.23 1.64 2.5[4] 800
4 76788.94 1/2 or 3/2 9400.96 5/2 40403.77 -5.36 5.6[4]555 if $J_{n}$=3/2. 700555 if $J_{n}$=3/2.
5 76895.40 3/2 9400.96 5/2 40403.77 101.1 3.0[3] [^6] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1. 4666Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1.
6 76445.96 1/2 or 3/2 9061.10 5/2 40063.91 -8.48 7.1[3]555 if $J_{n}$=3/2. 24555 if $J_{n}$=3/2.
7 76508.84 3/2 - 7/2 9061.10 5/2 40063.91 54.4 1.1[3] [^7] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1. 4777Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1.
8 68921.30 3/2 or 5/2 1521.90 5/2 32524.71 6.06 2.6[3] 28
9 75840.96 3/2 or 5/2 8460.35 5/2 39463.16 -12.73 5.9[3] 94
We are unable to match the measured energy levels of excited electronic states with energies $E\gtrsim\omega_{N}/2$ to the corresponding calculated wave functions, as the uncertainty in the calculated energy levels exceeds the spacing between the measured levels. An additional issue is the uncertainty in the electron angular momenta of the measured excited states. To address these challenges, we adopted the following approach.
In the first step, we identified the final electronic state $s$ and the intermediate electronic level $n$ that provide very small energy denominators $\Delta=E_{n}-E_{s}-\omega_{N}$ in Eq. (3). The smallest value of $\Delta$ is $-0.09$ cm-1. We also considered five additional cases with larger $\Delta$, where enhanced electronic bridge (EB) contributions may still be expected.
To estimate $\beta$, we evaluated the products of the matrix elements $x\equiv\langle s||T_{k}||n\rangle^{2}\langle n||D||t\rangle^{2}$ (see Eq. (3)) . The measured value of the final state level $s$, which has low energy, can be reliably matched with its corresponding theoretical energy level and wave function. However, we do not have definitive theoretical identifications for the initial state $t$ and the intermediate state $n$ (see the diagram in Fig.1).
For each final state $s$, we calculated 217 possible values of $x$, corresponding to different assignments of the experimental levels $t$ and $n$. The distribution of these $x$ values indicates that there is no strong configuration mixing. Such mixing would result in eigenstates made from chaotic superpositions of all configurations close in energy and smooth distribution of $x$ values. Instead, the distribution of $x$ consists of a small group of large values that dominate the average $\langle x\rangle$, and a broad distribution of smaller values. Thus, we cannot apply statistical theory developed for chaotic eigenstates [27, 28, 29, 30, 31].
The large matrix elements arise when the leading configurations in the initial and final states differ by a single orbital, allowing for a strong single-particle transition. The smaller matrix elements are generated by configuration mixing. To identify a representative value of $x$, we found the median value $x_{m}$ (in the middle of $x$ distribution). Note that $x_{m}$ is significantly smaller than the average $\langle x\rangle$.
The results of our calculations for several possible transitions are presented in Table 1. Our findings show that the probability of nuclear excitation can, in principle, be enhanced 5 million times. However, the most probable cases correspond to $\beta(x_{m})$, presented in the last column of the table. In these cases, the enhancement may reach five orders of magnitude when the magnetic dipole hyperfine interaction is allowed (for $J_{n}=3/2$). This enhancement can be achieved through proper selection of laser frequencies in a two-step atomic excitation process, followed by nuclear excitation.
IV Numerical calculation of the electronic bridge contribution to the decay of the nuclear excited state
Figure 2: An EB contribution to the decay rate of the isomeric nuclear state $m$. Index $i$ stands for the atomic ground state, $n$ stands for intermediate states, $f$ stands for final odd-parity states.
Table 2: The values of $\beta$ for the EB contribution to the decay rate of the isomeric nuclear state. The rate depends on the value of the unknown total angular momentum $J_{n}$ of the state which is closest in energy to the energy of nuclear transition $\omega_{N}$ = 67393.34 cm-1. Possible values of $J_{n}$ for $\Delta=E_{n}-\omega_{N}=14.73$ cm-1 are 3/2, 5/2 and 7/2. If $J_{n}=7/2$, then the magnetic dipole hfi does not contribute, contribution of the the electric quadrupole hfi is small ($\beta=2$) and the leading contribution comes from the next closest state with energy $E_{n}$ = 67509.63 cm-1, $\Delta=E_{n}-\omega_{N}=116.29$ and $J_{n}=3/2$ (spectra from Ref.  [23]). $x\equiv\langle i||T_{k}||n\rangle^{2}\langle n||D||f\rangle^{2}$; $\langle x\rangle$ is the average value, $x_{m}$ is the median value.
$J_{n}$ $E_{n}$ $\Delta$ $\beta(\langle x\rangle)$ $\beta(x_{m})$
[cm-1] [cm-1]
3/2 67378.61 14.73 3963 337
5/2 67378.61 14.73 5595 341
7/2 67378.61 14.73 2111Only electric quadrupole hfi contributes.
3/2 67509.63 116.29 63222Contribution from the next state $E_{n}$=67509.63  cm-1, in case of $J_{n}$=7/2 for $E_{n}$=67378.61 cm-1. 5
The electronic bridge contribution to the decay of the nuclear excited state is described by the diagram in Fig. 2. This process requires summing over all possible final electronic states $f$. The sum over the intermediate states $n$ is dominated by a single term with energy $E_{n}=67378.61$ cm-1, which results in a small energy denominator in the hyperfine interaction induced EB amplitude: $\Delta=E_{n}-\omega_{N}=14.73$ cm-1.
Unfortunately, the electron angular momentum for this level is unknown, so we have considered all three possible cases: $J_{n}=3/2$, 5/2, and 7/2. When $J_{n}=7/2$, only the electric quadrupole hfi contributes. A larger magnetic dipole hfi contribution arises from a higher level with $J_{n}=3/2$, at $E_{n}=67509.63$ cm-1, with $\Delta=E_{n}-\omega_{N}=116.29$ cm-1.
We cannot assign definite wave functions to the states $n$ and the high-energy final states $f$. Therefore, we calculate the matrix element product $x\equiv\langle i||T_{k}||n\rangle^{2}\langle n||D||f\rangle^{2}$ for the lowest state $f$ and approximately 30 possible assignments of states $n$. The average value $\langle x\rangle$ and the median value $x_{m}$ are used in the calculation of the enhancement factors $\beta$ The contribution of the higher $f$ states is estimated assuming the $\omega^{3}$ dependence of the individual terms (see Eq. (2) ). Approximately 50 experimental energy levels are used in the estimation. The saturation of the summation was clearly observed.
We present the enhancement factor $\beta$ for both the average value $\langle x\rangle$ and the median value $x_{m}$ in Table 2.
V Conclusion
We have identified several cases of resonance between electronic and nuclear excitations, indicating the potential for significant enhancement of nuclear excitation through the electronic bridge mechanism. This enhancement can be achieved by selecting appropriate laser frequencies in a two-step atomic excitation process, followed by nuclear excitation.
Due to uncertainties in assigning calculated wave functions to the experimentally measured energy levels near the nuclear excitation energy, it was not possible to provide precise enhancement factors. These factors may vary from a thousandfold to a millionfold enhancement. To address this uncertainty, we analyzed hundreds of possible level assignments and presented the average and median values of the enhancement factor $\beta$ for six selected laser frequency configurations exciting both nuclear and electronic transitions (see Table 1).
We also estimated the effect of electronic states on the decay lifetime of the nuclear excited state (see Table 2).
Accurate measurements of the magnetic moments and angular momenta of the Th II electronic states would help refine the theoretical identification of the measured levels and improve the accuracy of predictions for the electronic bridge mechanism.
This work was supported by the Australian Research Council Grant No. DP230101058.
Appendix A Method of Calculation
We use the same computational method as in our previous work on the ThIII ion[21]. The method is based on the relativistic Hartree-Fock (RHF) approach combined with the configuration interaction and single-double coupled cluster (CI+SD) method [32] to calculate the three-electron valence states of the Th II ion.
The calculations begin with the closed-shell ThV ion. The single-electron basis states for the valence electrons are generated in the field of the frozen core using the B-spline technique[33]. The SD equations are first solved for the core and subsequently for the valence states [32]. This procedure results in the creation of one-electron and two-electron correlation operators, $\Sigma_{1}$ and $\Sigma_{2}$, which are then applied in the CI calculations.
To compute the transition amplitudes, we employ the time-dependent Hartree-Fock method [34], which is equivalent to the well-known random-phase approximation (RPA). Solving the RPA equations for the valence states produces effective operators for the external field, which are used to evaluate matrix elements between the valence states.
It is important to note that the states of interest have excitation energies around 70,000 cm-1. For Th II, this requires calculating approximately one hundred states for each value of $J$ and parity. The density of states in this energy region is high, and the energy intervals between them are smaller than the uncertainty in the calculations.
Consequently, reliably matching the measured energy levels to the calculated levels and wave functions of the high-energy states is currently not feasible. To address this challenge, we employ the averaging procedure described earlier in the manuscript.