### v2

Abstract
The 8.4 eV transition in the 229Th nucleus is the basis for a high-precision nuclear clock with exceptional sensitivity to new physics effects. We have identified several cases in the Th+ ion where electronic excitations closely resonate with the nuclear excitation, with the smallest energy difference being $\Delta=-0.09$ cm-1. We investigate the electronic bridge process, in which nuclear excitation is induced via electronic transitions, and demonstrate that a proper selection of laser frequencies can lead to a dramatic enhancement of this effect. Additionally, we show that the interaction with electrons significantly shortens the lifetime of the nuclear excited state.
I Introduction
It was suggested in Ref.[1] that the 229Th nucleus has a low-energy metastable state. Ref.[2] proposed using the transition between the ground and metastable states as the basis for a nuclear clock. The precision of such a clock in a ”stretched” state, with maximal projections of both electronic and nuclear angular momenta, is expected to be very high because, in this state, the electron and nuclear variables are not entangled [3]. Indeed, in such a state, the nucleus is practically decoupled from external fields, such as those produced by blackbody radiation. This leads to a significant suppression of corresponding systematic effects.
On the other hand, the frequency of this transition is highly sensitive to effects of physics beyond the Standard Model, including violations of the Lorentz invariance and the Einstein equivalence principle [4], as well as variations in fundamental physical constants [5], such as the fine-structure constant $\alpha$, quark masses, and the QCD scale $\Lambda_{QCD}$. This sensitivity arises from the very small transition energy (8.4 eV), which is the result of strong cancellations between contributions from strong and electromagnetic nuclear interactions, typically at the MeV scale. This cancellation makes the transition frequency extremely sensitive to small changes in these contributions caused by variations in fundamental constants. Such variations may result from interactions with scalar or axion dark matter fields. Therefore, measuring these variations in clock experiments may provide an efficient method for searching for these fields [6, 7].
The measurement of the energy of the nuclear clock transition in 229Th has been an ongoing effort for many years (see, e.g., [8, 9, 10]). Recent advances in laser spectroscopy methods have led to a breakthrough in accuracy[11, 12, 13]. The transition frequency has been measured to be [12] $\omega_{N}=2,020,407,384,335(2)$ kHz [13] (67393 cm-1) in 229Th atoms embedded in a solid. Even higher accuracy is anticipated in future experiments with 229Th ions [3].
The amplitude of the nuclear M1 transition is suppressed by five orders of magnitude compared to typical atomic E1 transitions. However, it is possible to enhance this transition using the electronic bridge (EB) process, where electronic transitions induce nuclear transitions via magnetic interaction between electron and nucleus. The EB effect also increases the probability of decay of the nuclear excited state [14]. The high efficiency of excitation of the low-lying isomeric level in the 229Th nucleus by photons through the EB was predicted about 30 years ago in the works of Refs. [15, 16, 17]. More calculations have been performed for Th II, Th III, and Th IV ions, see e.g.  [18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Our previous EB calculations for the decay [19] and excitation [20] of the excited nuclear state in the Th II ion could not provide accurate results due to a large uncertainty in the measured value of the nuclear excitation energy and the absence of measurements of the positions of Th II electron energy levels near the nuclear excitation energy. According to theoretical expectations [28], the spacing between electron energy levels in this region is very small, making resonance with the nuclear excitation possible.
The aim of this paper is to propose an atomic excitation scheme that maximizes the EB effect and to estimate the magnitude of this effect. Additionally, we estimate the impact of the EB process on the decay of the nuclear excited state. Recent data on the Th II electron spectrum [31] (see Supplemental Material for the most complete data) allow us to identify several cases where a small energy denominator in the EB excitation formula leads to a very large resonance enhancement of the EB effect.
The error in the calculated level positions exceeds the very small spacing between the measured electron energy levels in highly excited states. To estimate the EB effects, we adopt a statistical approach (based on the calculations of many possible EB transition probabilities) due to the current lack of information needed to match measured energy levels with calculated electron wave functions. Accurate measurements of the magnetic moments and angular momenta of the Th II electron states would assist in the theoretical identification of the measured levels and improve the accuracy of EB predictions.
II Electronic bridge for nuclear excitation in Th II
We follow the method used in Refs. [20, 25] for the Th II anf Th III ions. The decay of any atomic state with energy $E_{n}>\omega_{N}$ may include nuclear excitation. Since the nuclear excitation energy $\omega_{N}=67393$ cm-1 lies outside the optical region, we consider a two-step excitation of the electronic states(see also [20]). In the first step, the atom is excited from the ground state (GS) to one of the intermediate states $t$ used in [20]. In the second step, the ion is further excited by a second laser with frequency which is chosen to satisfy the energy conservation condition for simultaneous nuclear excitation and excitation of the final electronic state, $\omega_{1}+\omega_{2}=\omega_{N}+E_{s}$ ($\omega_{1}=E_{t}$). Nuclear excitation is due to the hyperfine interaction between atomic electrons and the nucleus. Corresponding diagram is presented in Fig. 1. Its calculation involves a summation over all intermediate states $n$. However, in this analysis, we focus on the dominant contributions, which come from specific intermediate states $n$ and final states $s$ providing a very small energy denominator in Eq. (3). Note that the final state $s$ is not the ground state but a low-lying excited state satisfying selection rules for the E1 transition induced by magnetic interaction between electron and nucleus, from the state $t$ (see diagram in Fig. 1). Suggestion that EB excitation process with final electron in an excited state has higher probability was made in Ref.  [29].
We do not go to highly excited states $s$ to keep laser frequencies within optical region. For each final electron state $s$, we look for an intermediate state $n$ (using energy level data from Ref. [31]) which corresponds to the smallest energy denominator $\Delta$ in (3), $\Delta=E_{n}-E_{s}-\omega_{N}$, and use this state for estimation of the EB amplitude.
Figure 1: A diagram for the EB excitation process in Th II. It is assumed that the atom is in initial electronic state $t$; $n$ and $s$ are intermediate and final electronic states, $g$ ground nuclear state, $m$ is isomeric nuclear state.
The rate of an induced excitation from state $b$ to state $a$ $W_{ba}^{\rm in}$ can be calculated using the rate of a spontaneous transition $a\rightarrow b$ $W_{ab}$ [33]:
$W_{ba}^{\rm in}=W_{ab}\frac{4\pi^{3}c^{2}}{\omega^{3}}I_{\omega}.$ (1)
Here $I_{\omega}$ is the intensity of isotropic and unpolarized incident radiation. Index $b$ corresponds to nuclear ground state and electron excited state, index $a$ corresponds to nuclear excited state and electronic ground or low energy excited state. The rate of a spontaneous transition via electronic bridge process is given by [18]
$W_{ab}=\frac{4}{9}\left(\frac{\omega}{c}\right)^{3}\frac{|\langle I_{g}||M_{k}||I_{m}\rangle|^{2}}{(2I_{m}+1)(2J_{t}+1)}G_{2}^{(k)},$ (2)
Keeping in mind the relation (1), we assume that $\omega$ in (2) is the frequency of second excitation ($\omega\equiv\omega_{2}$). Factor $G_{2}$ in (2) depends on electrons only. It corresponds to upper line of Fig. 1. Assuming resonance situation and keeping only one strongly dominating term, we have
$G_{2}^{(k)}\approx\frac{1}{2J_{n}+1}\left[\frac{\langle s||T_{k}||n\rangle\langle n||D||t\rangle}{\omega_{ns}-\omega_{N}}\right]^{2}.$ (3)
Here $T_{k}$ is the electron part of the hyperfine interaction operator (magnetic dipole (M1) for $k=1$ and electric quadrupole (E2) for $k=2$), $D$ is the electric dipole operator (E1). States $n$ and $s$ are chosen to get close to a resonance, $\epsilon_{n}-\epsilon_{s}\equiv\omega_{ns}\approx\omega_{N}$.
It is convenient to present the results in terms of dimensionless ratios ($\beta$) of EB transition rates to the nuclear transition rates.
$\beta_{M1}=\Gamma^{(1)}_{\rm EB}/\Gamma_{N}(M1),\ \ \beta_{E2}=\Gamma^{(2)}_{\rm EB}/\Gamma_{N}(E2).$ (4)
Here $\Gamma^{(k)}$ is given by (2) ($\Gamma^{(k)}\equiv W_{ab}$). Both parameters, $\beta_{M1}$ and $\beta_{E2}$ can be expressed via $G2$, (Eq. (3)) [21]
$\beta_{M1}=\left(\frac{\omega}{\omega_{N}}\right)^{3}\frac{G_{2}^{(1)}}{3(2J_{s}+1)},$ (5)
$\beta_{E2}=\left(\frac{\omega}{\omega_{N}}\right)^{3}\frac{4G_{2}^{(2)}}{k_{n}^{2}(2J_{s}+1)}.$ (6)
The $M1$ and $E2$ contributions can be combined into one effective parameter $\tilde{\beta}$ using the known ratio of the widths of the nuclear $M1$ and $E2$ transitions, $\gamma=\Gamma_{\gamma}(E2)/\Gamma_{\gamma}(M1)=6.9\times 10^{-10}$, [34]. Then $\tilde{\beta}=\beta_{M1}(1+\rho)$, where $\rho=\gamma\beta_{E2}/\beta_{M1}$.
III Numerical calculations of the electronic bridge contribution to the nuclear excitation
Table 1: Versions of the EB process with resonant nuclear excitation. The notations $t$, $n$, and $s$ refer to the diagram in Fig. 1. It is assumed that $E_{t}$ = 36390.53 cm-1, $\omega_{2}=\omega_{N}+E_{s}-E_{t}$ is the frequency of second laser excitation. Final states $s$ are listed with their experimental energies ($E_{s}$) and calculated lifetimes ($\tau_{s}$). The energy denominator in Eq. (3) is given by $\Delta=E_{n}-E_{s}-\omega_{N}$. The third column lists all possible values of $J_{n}$ from the experimental data in Ref. [31], the last two columns present the calculated values of the EB enhancement factors $\beta$ for specific $J_{n}$. Here, $x\equiv\langle s||T_{k}||n\rangle^{2}\langle n||D||t\rangle^{2}$, and $\langle x\rangle$ represents the average value of $x$ over theoretical states that are candidates to represent the actual states $t$ and $n$. No averaging is performed over $s$ since it is a low-lying state that is easily identified in the calculations. The median value of $x$ is denoted by $x_{m}$. Numbers in square brackets indicate powers of ten.
$N$ $E_{n}$ [^1] Ref. [31], Supplemental Material. $J_{n}$ State $s$ $E_{s}$222 NIST database, [32]. $J_{s}$ $\tau_{s}$ $\omega_{2}$ $\Delta$ $\beta(\langle x\rangle)$ $\beta(x_{m})$
[cm-1] [cm-1] [cm-1] [cm-1]
1 73637.54 3/2 or 5/2 $6d^{2}7s$ 4P1/2 6244.29 1/2 7.8 s 37247.10 -0.09 5.5[6]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole interaction between electron and nucleus. 1.6[5]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole interaction between electron and nucleus.
2 73637.54 3/2 or 5/2 $6d^{2}7s$ 4P1/2 6244.29 1/2 37247.10 -0.09 1.2[4]444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole interaction between electron and nucleus contributes. 660444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole interaction between electron and nucleus contributes.
3 74396.40 3/2 or 5/2 $6d^{3}$ 3/2 7001.42 3/2 110 ms 38004.23 1.64 2.5[4] 800
4 76788.94 1/2 or 3/2 $6d^{3}$ 4F5/2 9400.96 5/2 1.5 ms 40403.77 -5.36 5.6[4]555 if $J_{n}$=3/2. 700555 if $J_{n}$=3/2.
5 76895.40 3/2 $6d^{3}$ 4F5/2 9400.96 5/2 40403.77 101.1 3.0[3] [^6] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1. 4666Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1.
6 76445.96 1/2 or 3/2 $6d^{2}7s$ 4P5/2 9061.10 5/2 0.7 ms 40063.91 -8.48 7.1[3]555 if $J_{n}$=3/2. 24555 if $J_{n}$=3/2.
7 76508.84 3/2 - 7/2 $6d^{2}7s$ 4P5/2 9061.10 5/2 40063.91 54.4 1.1[3] [^7] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1. 4777Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1.
8 68921.30 3/2 or 5/2 $6d^{2}7s$ 4F5/2 1521.90 5/2 12 s 32524.71 6.06 2.6[3] 28
9 75840.96 3/2 or 5/2 $6d^{3}$ 3/2 8460.35 5/2 2.6 ms 39463.16 -12.73 5.9[3] 94
We are unable to match the measured energy levels of excited electronic states with energies $E\gtrsim\omega_{N}/2$ to the corresponding calculated wave functions, as the uncertainty in the calculated energy levels exceeds the spacing between the measured levels. An additional issue is the uncertainty in the electron angular momenta of the measured excited states. To address these challenges, we adopted the following approach.
In the first step, we identified the final electronic state $s$ and the intermediate electronic level $n$ that provide very small energy denominators $\Delta=E_{n}-E_{s}-\omega_{N}$ in Eq. (3). The smallest value of $\Delta$ is $-0.09$ cm-1. We also considered five additional cases with larger $\Delta$, where enhanced electronic bridge contributions may still be expected.
To estimate $\beta$, we evaluated the products of the matrix elements $x\equiv\langle s||T_{k}||n\rangle^{2}\langle n||D||t\rangle^{2}$ (see Eq. (3)) . The measured value of the final state level $s$, which has low energy, can be reliably matched with its corresponding theoretical energy level and wave function. However, we do not have definitive theoretical identifications for the initial state $t$ and the intermediate state $n$ (see the diagram in Fig.1).
For each final state $s$, we calculated 217 possible values of $x$, corresponding to different assignments of the experimental levels $t$ and $n$. The distribution of these $x$ values indicates that there is no strong configuration mixing. Such mixing would result in eigenstates made from chaotic superpositions of all configurations close in energy and smooth distribution of $x$ values. Instead, the distribution of $x$ consists of a small group of large values that dominate the average $\langle x\rangle$, and a broad distribution of smaller values. A similar conclusion was reached in Ref. [35]. Thus, we cannot apply statistical theory developed for chaotic eigenstates [36, 37, 38, 39, 40].
The large matrix elements arise when the leading configurations in the initial and final states differ by a single orbital, allowing for a strong single-particle transition. The smaller matrix elements are generated by configuration mixing. To identify a representative value of $x$, we found the median value $x_{m}$ (in the middle of $x$ distribution). Note that $x_{m}$ is significantly smaller than the average $\langle x\rangle$.
The results of our calculations for several possible transitions are presented in Table 1. Our findings show that the probability of nuclear excitation can, in principle, be enhanced 5 million times. However, the most probable cases correspond to $\beta(x_{m})$, presented in the last column of the table. In these cases, the EB enhancement factor $\beta$ may reach five orders of magnitude when the magnetic dipole hyperfine interaction is allowed (for $J_{n}=3/2$). This enhancement can be achieved through proper selection of laser frequencies in a two-step atomic excitation process, followed by nuclear excitation.
However, as it was noted in Ref. [27], a very large value of EB enhancement factor $\beta$ does not necessarily lead to the $\beta$ times enhancement of the excitation rate. The point is that the photon capture resonance cross section contains total width in denominator, which may be dominated by the radiative width $\Gamma_{\gamma}$, if it is very strongly enhanced by the EB process ($\Gamma_{\gamma}=(1+\beta)\Gamma_{N}$, where $\Gamma_{N}\approx 5\times 10^{-4}$ Hz is the the decay width of the nuclear isomeric state in bare nucleus - see e.g [42]). This effect can be qualitatively described by the factor, which appears in the resonance photon capture cross section
$\sigma\sim\lambda^{2}\frac{\Gamma_{\gamma}\Gamma}{(\omega-\omega_{0})^{2}+\Gamma^{2}/4},$ (7)
where $\lambda$ is the photon wavelength, $\Gamma=\Gamma_{\gamma}+\Gamma_{Ni}+\Gamma_{s}+\Delta_{l}$, $\Gamma_{Ni}$ is width of the nuclear transition in 229mTh+ , $\Gamma_{s}$ is the width of the final electron state and $\Delta_{l}$ is the width of the energy distribution of the photons in the pumping laser beam (or width of the frequency comb teeth). There may be other line broadening mechanisms.
In the resonance $\sigma\propto\beta\Gamma_{N}/(\beta\Gamma_{N}+\Gamma_{Ni}+\Gamma_{s}+\Delta_{l})$. We see that in the limit $\beta\to\infty$ the enhancement in the center of the resonance disappears.
However, in a realistic situation there is $\beta$ times enhancement of the cross section. Consider different contributions to the total width $\Gamma$. Specifically, for the final electron states 4P1/2 with the energy 6244.29 cm-1 and 5F5/2 with the energy 1521.90 cm-1 , the radiative width $\Gamma_{s}$ is very small, 0.1 Hz, since it is produced by low frequency M1 and E2 transitions. For other final states in the Table 1, E1 transitions are possible, and the width is from 10 to 1000 Hz. Note that all lifetimes are strongly affected by the configuration mixing.
There is an experimental indication that the lifetime of the nuclear isomer in 229mTh+ with electrons in their ground state is smaller than 10 ms [41] (see discussion in the next section), i.e. in 229mTh+ the natural decay width $\Gamma_{Ni}>$ 100 Hz. When electrons are in exited state, $\Gamma_{Ni}$ would hardly be smaller. If this is the case, there is an enhancement of the cross section at least up to $\beta=2\times 10^{5}$.
Moreover, during the frequency scanning procedure we are not exactly in the resonance, i.e. $(\omega-\omega_{0})^{2}>\Gamma^{2}/4$. The width of the frequency comb teeth $\Delta_{l}$ may significantly exceed natural width $\Gamma$. Finally, the scanning time is reduced inversely proportional to the width, so effectively one should consider the product of the cross section times the total width, so $\beta$ in the denominator cancels out. In this case the EB enhancement factor $\beta$ is important for any large $\beta$.
IV Numerical calculation of the electronic bridge contribution to the decay of the nuclear excited state
Figure 2: An EB contribution to the decay rate of the isomeric nuclear state $m$. Index $i$ stands for the atomic ground state, $n$ stands for intermediate states, $f$ stands for final odd-parity states.
Table 2: Calculations of the coefficient $R$ (8) at different assumptions for experimental level at $E_{n}$=67378 cm-1 ($J_{n}$=3/2,5/2 or 7/2). $N$ is serial number of the state in the calculations for given $J_{n}$ and parity; $\Delta E_{n}$ is the shift in the energy of state $n$ to bring it to experimental energy of 67378 cm-1. Numbers in square brackets stand for powers of ten.
$N$ $E_{n}$ cm-1 $\Delta E_{n}$ Partial $J_{f}$ contributions Sum
$J_{n}$=3/2; $J_{f}:$ 1/2 3/2 5/2
68 67601 222 8.524[1] 1.414[2] 8.824[1] 3.149[2]
69 67914 536 1.908[3] 3.929[3] 1.695[3] 7.532[3]
70 68174 796 4.276[2] 5.208[2] 2.036[2] 1.152[3]
71 68568 1189 6.270[3] 3.274[3] 1.933[3] 1.148[4]
72 68866 1488 2.554[2] 2.161[2] 2.261[2] 6.976[2]
73 69131 1752 6.223[2] 6.795[2] 1.279[3] 2.580[3]
74 69391 2013 1.762[3] 1.398[3] 2.280[3] 5.440[3]
75 69500 2121 1.548[3] 4.602[3] 4.080[3] 1.023[4]
76 69719 2340 8.695[3] 6.655[3] 1.121[4] 2.655[4]
77 69874 2496 5.716[2] 5.475[2] 6.200[2] 1.739[3]
78 70221 2842 5.813[2] 2.826[3] 1.475[3] 4.882[3]
79 70583 3204 2.090[3] 6.944[3] 8.203[3] 1.724[4]
80 71018 3639 1.665[3] 1.943[3] 3.999[3] 7.607[3]
81 71104 3725 2.949[2] 3.232[2] 5.333[2] 1.151[3]
82 71270 3892 3.868[3] 4.111[3] 9.680[3] 1.766[4]
83 71528 4149 1.113[2] 6.798[1] 1.302[2] 3.095[2]
84 71816 4437 1.108[3] 7.471[2] 4.730[3] 6.586[3]
85 72150 4771 2.141[3] 4.601[3] 1.353[4] 2.028[4]
$J_{n}$=5/2: $J_{f}:$ 3/2 5/2 7/2
78 67493 115 1.624[4] 2.049[4] 4.874[3] 4.160[4]
79 67915 536 9.062[2] 4.504[2] 2.893[2] 1.646[3]
80 68350 971 2.976[4] 1.724[4] 2.396[4] 7.096[4]
81 68531 1153 3.824[4] 5.077[4] 4.200[4] 1.310[5]
82 69006 1628 5.328[2] 1.447[3] 7.728[2] 2.752[3]
83 69196 1818 1.458[4] 3.942[4] 3.654[4] 9.054[4]
84 69309 1930 6.857[3] 5.275[3] 2.767[3] 1.490[4]
85 69352 1973 9.055[2] 1.242[3] 6.894[2] 2.837[3]
86 69640 2261 8.408[3] 2.196[4] 1.227[4] 4.264[4]
87 69817 2439 6.820[3] 2.231[4] 1.243[4] 4.155[4]
88 69918 2539 3.727[2] 8.864[2] 6.485[2] 1.908[3]
89 69974 2596 6.553[2] 1.908[3] 2.102[3] 4.665[3]
90 70222 2844 1.637[3] 3.367[3] 2.770[3] 7.774[3]
91 70475 3096 5.853[3] 6.302[3] 3.767[3] 1.592[4]
92 70510 3131 9.887[3] 9.563[3] 5.978[3] 2.543[4]
93 70823 3444 9.029[1] 1.072[2] 6.852[1] 2.660[2]
94 70924 3546 8.687[2] 9.090[2] 5.178[2] 2.296[3]
95 71035 3656 2.439[2] 2.602[2] 1.229[2] 6.270[2]
The electronic bridge contribution to the decay of the nuclear excited state is described by the diagram in Fig. 2. This process requires summing over all possible final electronic states $f$. The sum over the intermediate states $n$ is dominated by a single term with energy $E_{n}=67378.61$ cm-1 [31], which results in a small energy denominator in EB amplitude induced by magnetic dipole or electric quadrupole interaction with the nucleus: $\Delta=E_{n}-\omega_{N}=14.73$ cm-1. Very little is known about this state, only energy and a range of possible values of the total angular momentum ($J_{n}$=3/2, 5/2 or 7/2). Therefore, we consider a number of possible candidates (see below).
As before, it is convenient to present the results in terms of dimensionless ratios of EB transition rates to the nuclear transition rate. Note, that summation over final states $f$ leads to summation of terms similar to $\beta_{M1}$ (5) and $\beta_{E2}$ (6) with different values of $\omega$: $\omega\equiv\omega_{f}=\omega_{N}-E_{f}-E_{i}$. Therefore, we introduce new variable $R$
$R=\sum_{f}\beta(\omega_{f}).$ (8)
In the paper  [41] they study internal conversion signal which should appear after collision of the 229mTh isomer ions with a solid surface. They see clear signals from Th III and Th IV ions but do not observe any signal from Th II ions. The authors of Ref.  [41] suggested a possible explanation: lifetime ($\tau$) of the 229mTh isomer is very small, $\tau<10~{}{\rm ms}$, so Th II ions do not reach the surface.
The half-life of the isomeric nuclear state in Th IV is practically not affected by the electronic bridge mechanism [18]. This half-life in Th IV was measured to be $T_{1/2}=1400^{+600}_{-300}$ [42], corresponding to the lifetime $\tau=2020^{+866}_{-433}$. This lifetime is in a reasonable agreement with the values obtained in the solid state experiments [11, 12, 13, 43, 47, 44, 45, 46, 48] after introducing the solid state corrections. To obtain $\tau<10~{}{\rm ms}$, one needs very large $R>2\times 10^{5}$. This small value of the lifetime constitutes the so called thorium puzzle. Possible solutions of this puzzle (very large value of EB coefficient, $R=56500$, collisional broadening) were discussed in Refs.  [35, 29, 30].
We study this problem in present work using new experimental data on Th+ energy levels [31]. Strong enhancement of the decay rate due to above mentioned small energy denominator ($\Delta=14.73$ cm-1) may potentially explain the puzzle. We perform the calculations summing over large number of intermediate states $n$ around $E_{n}\sim\omega_{N}$ (about 20 states for each $J_{n}$) and all final states with $E_{f}<\omega_{N}$ (about 70 states for each $J_{f}$).
Comparison with experiment for excitation energies between 0 and 60000 cm-1 shows that when approaching the upper limit 60000 cm-1, calculated energies tend to be about 3000 cm-1 higher than experiment. Since we need states with energies $E_{n}\sim\omega_{N}$ = 67393.34 cm-1, it does make sense to shift calculated energies down to make more realistic energy denominators.
[…]
The largest values of $R$ in the Table, $R=1.3\times 10^{5}$, corresponds to the assumption that the experimental energy $E_{n}$=67378.61 cm-1 match up with the calculated state number 81 with the calculated energy 68531 cm-1 and total angular momentum $J_{n}=5/2$. This maximal calculated value of $R$, $R=1.3\times 10^{5}$, is close to the experimental lower bound, $R>2\times 10^{5}$. Therefore, we cannot entirely exclude this scenario as a potential solution to the thorium puzzle: an unexpectedly large enhancement of the nuclear decay rate via the electronic bridge mechanism [35, 29, 30].
We have also performed calculations for many levels located below nuclear frequency $\omega_{N}$ = 67393.34 cm-1, but they do not produce any qualitative changes to the picture.
V Conclusion
We have identified several cases of resonance between electronic and nuclear excitations, indicating the potential for significant enhancement of nuclear excitation through the electronic bridge mechanism. This enhancement can be achieved by selecting appropriate laser frequencies in a two-step atomic excitation process, followed by nuclear excitation.
Due to uncertainties in assigning calculated wave functions to the experimentally measured energy levels near the nuclear excitation energy, it was not possible to provide precise enhancement factors. These factors may vary from a thousandfold to a millionfold enhancement. To address this uncertainty, we analyzed hundreds of possible level assignments and presented the average and median values of the enhancement factor $\beta$ for six selected laser frequency configurations exciting both nuclear and electronic transitions (see Table 1).
We also estimated the effect of electronic states on the decay lifetime of the nuclear excited state (see Table 2). We identified one possible assignment of calculated wave functions and energy levels to observed electron energy levels that is consistent with the current experimental limit on the lifetime of ${}^{229\text{m}}$h. However, to explain the unexpectedly short lifetime of the ${}^{229\text{m}}$h state in Th II, it is important to experimentally investigate the possible existence of additional electronic levels near the energy of the nuclear isomeric state.
Accurate measurements of the magnetic moments and angular momenta of the Th II electronic states would help to refine the theoretical identification of the measured levels and improve the accuracy of predictions for the electronic bridge mechanism.
This work was supported by the Australian Research Council Grant No. DP230101058. We are grateful to Feodor Karpeshin for valuable discussions and references.
Appendix A Method of Calculation
We use the same computational method as in our previous work on the Th III ion [25]. The method is based on the relativistic Hartree-Fock (RHF) approach combined with the configuration interaction and single-double coupled cluster (CI+SD) method [49] to calculate the three-electron valence states of the Th II ion.
The calculations begin with the closed-shell ThV ion. The single-electron basis states for the valence electrons are generated in the field of the frozen core using the B-spline technique [50]. The SD equations are first solved for the core and subsequently for the valence states [49]. This procedure results in the creation of one-electron and two-electron correlation operators, $\Sigma_{1}$ and $\Sigma_{2}$, which are then applied in the CI calculations.
To compute the transition amplitudes, we employ the time-dependent Hartree-Fock method [51], which is equivalent to the well-known random-phase approximation (RPA). Solving the RPA equations for the valence states produces effective operators for the external field, which are used to evaluate matrix elements between the valence states.
It is important to note that the states of interest have excitation energies around 70,000 cm-1. For Th II, this requires calculating approximately one hundred states for each value of $J$ and parity. The density of states in this energy region is high, and the energy intervals between them are smaller than the uncertainty in the calculations.
Consequently, reliably matching the measured energy levels to the calculated levels and wave functions of the high-energy states is currently not feasible. To address this challenge, we employ the averaging procedure described earlier in the manuscript.